OpenStudy (anonymous):

divide (x^3-y^3) by (x-y)

Directrix (directrix):

First, factor (x^3-y^3). One of the factors of (x^3-y^3) will divide out with the (x-y)

OpenStudy (anonymous):

can u show me the solution @Directrix ? thanks so much for help..

Directrix (directrix):

I am attaching a chart about the factoring pattern for the sum and difference of two cubes. Look at it and then I will help you with the solution.

OpenStudy (anonymous):

thank you...

Directrix (directrix):

(x^3-y^3) = (x - y) ( x^2 + xy + y^2)

Directrix (directrix):

[ (x^3-y^3) ]/ (x-y) = [(x - y) ( x^2 + xy + y^2)] / (x - y) = x^2 + xy + y^2 --> This is the quotient, the answer. The (x-y) factors in the numerator and the denominator divided out. On the factoring of x^3 - y^3, that is just one of those things you use until you know and recognize it.

Directrix (directrix):

Questions?

OpenStudy (anonymous):

ur great.. thanks for the big help... :)